# SLIM: Scalable and Lightweight LiDAR Mapping in Urban Environments

### v1

@@ V-B2 Efficient NFR
Figure 7: Visualization for the Hessian matrix before marginalization. Noted that $\mathbf{H}_{ll}$ is block-diagonal and all of other blocks are sparse matrix in general. The dimension of $\mathbf{H}_{ll}$ is much larger than those of $\mathbf{H}_{nn}$ and $\mathbf{H}_{mm}$. Therefore, in the implementation, we only store its diagonal blocks, while all other matrices are saved using a sparse matrix structure.
In our implementation, we strategically arrange the Hessian blocks, positioning all landmark-related blocks in the top left corner, marginalized blocks (frames) in the bottom right corner, and retained blocks in the center. This arrangement partitions the full Hessian matrix into nine distinct blocks, enabling us to reformulate the original problem presented in Equation (19) as follows:
@@
As depicted in Figure 6, $\mathbf{H}_{ln}$, $\mathbf{H}_{nn}$, $\mathbf{V}_{l}$, $\mathbf{V}_{n}$ are sparse matrices, and $\mathbf{H}_{ll}$ is a block-diagonal matrix. Figure 7 presents a graphical understanding of the nine blocks.
@@
Finally, within the each single Jacobian block $\mathbf{J}_{r_{k}}$ and the essential blocks of $\mathbf{\Sigma}^{m}_{rr}$, i.e. $\mathbf{H}_{o}^{-1}$, all the recovered information matrices $\left\{\mathbf{\Lambda}_{r_{k}}\right\}$ could be calculated with Equation (21) in parallel. So far, we have introduced the construction of all recovered observations after marginalization, resulting in a sparse factor graph that is very close to the dense factor graph. The experimental section will demonstrate the necessity of the map-centric NFR for long-term mapping.
@@ VI Experiments
We conduct real-world experiments to evaluate the performance of the proposed SLIM system. We first introduce the experimental setup in Section VI-A. Subsequently, the following tests are designed to validate the four capabilities of SLIM and to address the motivations outlined in Section I:
@@
    Lightweightness. In Section VI-C, we compare the requirements for storage space with other map representations.
@@
    Scalability. The designed map-centric NFR is validated in Section VI-D individually with multi-session maps.
@@
    Localizability. Section VI-E demonstrates that the maps generated from SLIM can be re-used for online robot localization.
@@ VI-A Experimental Set-up
Figure 8: LiDAR sensors and the LiDAR scans used in the experimental sections. The LiDAR point clouds vary due to the different specifications of LiDAR sensors. The front-end map vectorization module can still parametrize the point clouds to lines and planes, thus guaranteeing the following map refinement and maintenance modules.
@@
((b)) HeLiPR Roundabout
Figure 9: The multi-session datasets in NCLT and Roundabout, HeLiPR. We utilize the regions, a type of large-scale submaps, to better manage urban mapping. The zoomed-in views provide trajectories of multi-session robot travels with different colors and heights, and also the frames in the regions.
Three real-world datasets are employed for comprehensive validation and performance evaluation: KITTI [91], NCLT [92], and HeLiPR [9]. The KITTI dataset is collected by a vehicle equipped with a Velodyne HDL-64E. We use Sequence 00-10 for experimental validation, although some sequences are not entirely in urban environments. We employ KISS-ICP [14] to obtain LiDAR odometry as prior poses. Each sequence is partitioned into multiple sessions with overlapping regions before proceeding with our map merging pipeline. NCLT is also a well-known dataset suitable for the evaluation of multi-session mapping [69]. Multiple sessions are collected from a mobile platform equipped with a Velodyne HDL-32E. Due to the rotational motion of the platform, we use FAST-LIO2 [4] for point cloud undistortion and LiDAR odometry generation. We selected 6 regions as our benchmark, with each region containing between 25 and 60 sessions, meaning the robot travels to the location more than 25 times.
@@
As mentioned above, we partition the large-scale urban areas into regions for evaluation. Essentially, the region represents a type of submap for long-term management, which differs from the concept of session. More specifically, the former is in the spatial domain, while the latter is in the temporal domain. We present the visualization results of the multi-session data and the regions in Figure 9 for a better illustration. Figure 10 also shows the mapping results in different regions.
@@ VI-B Mapping Accuracy
((a)) NCLT Region 02
((b)) NCLT Region 04
((c)) HeLiPR DCC Region 05
((d)) HeLiPR KAIST Region 00
((e)) HeLiPR Town Region 02
((f)) HeLiPR Town Region 05
@@
This section demonstrates that SLIM can provide a globally consistent map from multi-session data. We first compare SLIM with other advanced mapping systems on KITTI, using trajectory accuracy for quantitative evaluation. Then, we demonstrate long-term map merging and consistent mapping with incremental map sessions on the HeLiPR dataset.
@@ VI-B1 Mapping Accuracy on KITTI Dataset
Regarding the evaluation metrics, we compute the absolute translation error between the keyframes in our global map and the corresponding frames in the ground truth trajectory. The benchmark results are summarized in Table I, with partial results obtained from BALM2 [12]. The overall results indicate that both our method and the state-of-the-art BALM2 [12] achieve the best performance in mapping accuracy. Notably, BALM2 utilizes dense point clouds for loop closure detection and pose optimization, while SLIM employs different representations. Additionally, we observe that SLIM performs poorly on Sequences 01 and 09 due to insufficient roadside features, such as streetlights and trees. As a sparse and lightweight LiDAR mapping system, the proposed SLIM still achieves SOTA-level mapping accuracy while maintaining map merge capability.
@@ VI-B2 Incremental Mapping on HeLiPR Dataset
((a)) Roundabout Region 00
((b)) Roundabout Region 01
((c)) Roundabout Region 02
((d)) Roundabout Region 03
((e)) Town Region 00
((f)) Town Region 01
((g)) Town Region 02
((h)) Town Region 03
((i)) DCC Region 00
((j)) DCC Region 01
((k)) DCC Region 02
((l)) DCC Region 03
((m)) KAIST Region 00
((n)) KAIST Region 01
((o)) KAIST Region 02
((p)) KAIST Region 03
Figure 11: Quantitative results of mapping accuracy on HeLiPR. The evaluation metric is the RMSE of absolute trajectory error. The overall results indicate that: (1) the accuracy is improved with the PGO followed by the LiDAR BA, displayed with purple and cyan colors, respectively; (2) the final optimized error after BA is bounded with increased map sessions.
Compared to the KITTI dataset, the recently published HeLiPR dataset is more suitable for validating long-term mapping. This is due to the inter-LiDAR sequences in HeLiPR, which naturally verify map merging and refinement as sessions increase at the same location. Specifically, the HeLiPR dataset encompasses sequences collected at four different scenarios: DCC, KAIST, Roundabout, and Town, with three sequences at each scenario. To obtain multiple submaps with overlapping regions, we selected 32 regions as our benchmark, with each region containing around 10 sessions. The SLIM system sequentially merges multiple submaps onto the base map, achieving incremental mapping from multi-session data. As depicted in Algorithm 1, SLIM conducts PGO and BA when a new session is merged, allowing for evaluation of the merged trajectory accuracy. We also compute the trajectory accuracy of the current submap to be merged. The heights of ground truth trajectories are not stable in some sequences of HeLiPR, so we evaluate trajectory accuracy along the $x$ and $y$ only.
((a)) DCC Region 03
((b)) Roundabout Region 00
((c)) Town Region 00
((d)) Town Region 05
Figure 12: Visualization of point cloud maps in HeLiPR regions. The point cloud maps are built by accumulating LiDAR points on robot poses. The poses on the left side are obtained before the LiDAR bundle adjustment, i.e., using the LiDAR odometry and pose graph optimization. After applying the LiDAR bundle adjustment, the poses on the right have fewer drifts, resulting in high-precision maps in zoomed-in views. This figure demonstrates the effectiveness of the proposed LiDAR bundle adjustment on line and plane representations.
We present the experimental results of 20 sequences quantitatively in Figure 11. The SLIM system can maintain the consistency of the global map as new maps are added, even if the new maps are aligned with poor prior poses. Notably, the error of BA is generally less than PGO and the original LiDAR odometry, demonstrating the effectiveness of the proposed map refinement in Section IV-C. We also present qualitative results by applying our proposed BA in Figure 12, verifying the necessity of using BA for high-consistency mapping. Additionally, we observe unstable results after applying PGO due to low-quality relative poses from block registration when prior poses of LiDAR odometry are not accurate enough. In summary, given multi-session maps, the map refinement modules can effectively reduce pose errors, i.e., improve mapping accuracy, by utilizing map merging and refinement with lightweight lines and planes.
One important capability of our framework is map merging, as described in Section IV-B, which is harder to evaluate quantitatively compared to mapping accuracy. All the input submaps in this section are within their own coordinates when generated. The map merge module can align them into one global frame, resulting in the error reduction shown in Figure 11. The video in the supplementary materials also presents the map merge process to facilitate understanding.
VI-C Memory Efficiency
TABLE II: Memory Consumption (Mega Byte, MB) on KITTI
Sequence (Length) 00 (3.7 Km) 05 (2.2 Km) 08 (3.2 Km)
Raw point cloud 13624.2 8284.7 12214.1
Point cloud ($r$=0.1m) 1773.40 1236.75 1254.92
Point cloud ($r$=0.3m) 23.48 16.77 18.56
Surfel map [97] 887.7 512.6 835.7
Mesh map [98] 2032.9 1317.4 1894.1
VDB TSDF map [99] 748.1 434.6 958.6
SHINE map [55, 57] 160.6 114.2 189.9
PIN SLAM [20] 102.1 66.3 138.8
SLIM 12.9 7.2 9.2
SLIM (L) 0.5 0.3 0.4
((a)) NCLT Region 00
((b)) NCLT Region 01
((c)) NCLT Region 02
((d)) NCLT Region 03
Figure 13: Quantitative results of map consumption of SLIM and the time cost for map refinement. The results on NCLT indicate that (1) the provided map is lightweight and only uses low storage consumption; (2) the consumption and time cost remain stable as the number of sessions increases.
Urban environments cover large spaces and various scenarios, so memory efficiency is a major concern for mobile robots. In this section, we evaluate the memory efficiency of SLIM compared to several conventional mapping systems in terms of storage size, i.e., the memory consumption for a robot. The SLIM system utilizes lines and planes as basic front-end representations. The stored map can also contain graph structures that encode poses and co-visible information. If the map is solely designed for localization without such structures, we denote it with an (L) label for clarity.
The memory consumption of different maps is first evaluated on the KITTI dataset. The comparisons include several conventional maps, such as sampled point clouds (with $r$ as the sampling radius) and implicit neural maps [55, 20]. Partial results are from previous studies [8, 20], and all are summarized in Table II. Our lightweight map requires significantly less storage than other mapping systems, demonstrating its superiority in memory efficiency for urban environments. Specifically, SLIM (L) only needs less than 0.5 Megabyte (MB) to cover a travel length of over 3 kilometers (km), with a density of approximately 130 KB/km. If the full SLIM map is needed for updates, the system uses less than 15 MB per sequence to store on disks.
The memory consumption on NCLT and HeLiPR is also presented in Figure 13 and Figure 14, respectively. In Figure 13, each region requires less than 1 MB to store a full map structure. Similar results are found in Figure 14. In summary, the built maps require low storage consumption on these two datasets, consistent with the maps generated from KITTI. High memory efficiency can reduce data transmission load in multi-robot systems for distributed localization and mapping, and long-term crowd-sourced urban mapping.
VI-D Scalability for Long-term Mapping
Sections VI-B and VI-C demonstrate that the SLIM system provides not only accurate but also memory-efficient maps in urban environments. The map merge capability is also verified in Section VI-B. As discussed in Section I, one motivation for SLIM is scalability, which is crucial for long-term mapping to maintain both consistency and manageable map sizes as map sessions increase. Furthermore, it is important that the time cost for map refinement, particularly for high-dimensional LiDAR BA, remains controllable.
We verify the scalability of SLIM on NCLT and HeLiPR, two classical long-term datasets. For NCLT, Figure 13 shows that both map consumption and time cost remain within certain limits even with 50 sessions in the same region. This satisfying performance is primarily attributed to the map-centric NFR module. To verify this, we disable the NFR and conduct tests on HeLiPR. In Figure 14, if NFR is disabled (w/o), we observe that map optimization time exhibits near-exponential growth, while map storage consumption increases linearly with the number of sessions. Overall, the full system maintains mapping costs on both NCLT and HeLiPR, thus verifying the scalability of the proposed SLIM system. Note that we do not provide experimental results on NCLT without NFR, as disabling NFR would extend submap merging time to tens of minutes or even hours.
One might argue that keyframe sparsification or similar sampling techniques could maintain a sparse map structure, i.e., the factor graph topology, for long-term mapping, making the NFR unnecessary due to its time costs in solving Equation (16). To verify this, we applied keyframe sparsification in our method while maintaining the prior information described in Section IV-A1, without solving the NFR problem. Specifically, we define this prior information as $\sqrt{\mathbf{\Lambda}_{r_{k}}}=\mathbf{I}\cdot\sqrt{N_{r_{k}}}/\sigma$, considering the sum of points in the original observations, as the number of points directly reflects their weights. Figure 15 shows that mapping accuracy drops compared to the full NFR module. These results demonstrate that NFR plays a vital role in both scalability and accuracy for long-term mapping.
By employing a minimum spanning tree to generate the topology of keyframe connections, we ensure that this topology remains as compact as possible. The absence of connections might result in subsequent map optimization failures, whereas adding any other connection is redundant. Figure 16 demonstrates the effectiveness of our keyframe sparsification approach, ensuring comprehensive connectivity among all keyframes while minimizing unnecessary information storage. An additional issue in topology reconstruction is whether the system can preserve mapping accuracy when a landmark is retained with only a single observation on the keyframe, i.e., each landmark is connected only to the nearest keyframe pose, as described in Section V-B. Our prior experiments in Figure 11 show that the system consistently maintained high trajectory accuracy following each BA step, resulting in high consistency in each refined base map. The optimized results of the LiDAR BA remain unaffected by the precision of the multi-session input submaps. These results indicate that the proposed map-centric NFR can provide a reasonable information matrix to ensure the stability of mapping refinement, even when each landmark is connected to one keyframe pose after topology reconstruction.

### v2

@@ V-B2 Efficient NFR
Figure 7: Visualization for the Hessian matrix before marginalization. Noted that $\mathbf{H}_{ll}$ is block-diagonal and all of other blocks are sparse matrices in general. $\mathbf{H}_{ll}$ is much larger in dimensions than $\mathbf{H}_{nn}$ and $\mathbf{H}_{mm}$. Therefore, in the implementation, we only store its diagonal blocks, while all other matrices are saved using a sparse matrices.
In our implementation, we strategically arrange the Hessian blocks, positioning all landmark-related blocks in the top left corner, marginalized blocks (frames) in the bottom right corner, and retained blocks in the center. This arrangement partitions the full Hessian matrix into nine distinct blocks, enabling us to reformulate the original problem presented in Equation (21) as Equations (26), (27) and (28). As depicted in Figure 6, $\mathbf{H}_{ln}$, $\mathbf{H}_{nn}$, $\mathbf{V}_{l}$, $\mathbf{V}_{n}$ are sparse matrices and $\mathbf{H}_{ll}$ is a block-diagonal matrix. Figure 7 presents a graphical understanding of the nine blocks.
@@
Finally, within the each single Jacobian block $\mathbf{J}_{r_{k}}$ and the essential blocks of $\mathbf{\Sigma}^{m}_{rr}$, i.e. $\mathbf{H}_{o}^{-1}$, all the recovered information matrices $\left\{\mathbf{\Lambda}_{r_{k}}\right\}$ could be calculated with Equation (23) in parallel. So far, we have introduced the construction of all recovered observations after marginalization, resulting in a sparse factor graph that is very close to the dense factor graph. The experimental section will demonstrate the necessity of map-centric marginalization and the effectiveness of our proposed NFR.
@@ VI Experiments
We conduct real-world experiments to evaluate the performance of the proposed SLIM system. We first introduce the experimental setup in Section VI-A. The following tests are designed to validate the four capabilities of SLIM and to address the motivations outlined in Section I:
@@
    Lightweightness. In Section VI-C, we compare the requirements for storage space (memory consumption) with other map representations.
@@
    Scalability. The designed marginalization is validated in Section VI-D individually with multi-session maps.
@@
    Localizability. Section VI-E demonstrates that the maps from SLIM can be re-used for online robot localization.
@@ VI-A Experimental Set-up
Figure 8: LiDAR sensors and scans in KITTI, NCLT and HeLiPR. The origin of each scan, corresponding to the sensor frame, is positioned at the center of the LiDAR point clouds. The properties of point clouds, such as density, vary with different sensor configurations. The front-end map vectorization module can still parametrize the point clouds to lines and planes, ensuring the functionality of the subsequent map refinement and maintenance modules.
@@
Figure 9: The multi-session datasets in NCLT. We employ the regions, a type of large-scale submaps, to better manage urban mapping. The zoomed-in views in the red, green and blue boxes provide trajectories of multi-session robot travels, visualized with varying heights and colors. The yellow boxes provide zoomed keyframe views of poses on the x-y plane.
Three real-world datasets are employed for comprehensive validation and quantitative evaluation: KITTI [89], NCLT [90], and HeLiPR [9]. The KITTI dataset is collected by a vehicle equipped with a Velodyne HDL-64E. We use Sequence 00-10 for experimental validation, although some sequences are not entirely in urban environments. We employ KISS-ICP [14] to obtain LiDAR odometry as prior poses. Each sequence is partitioned into multiple sessions with overlapping regions before proceeding with our map merging pipeline. NCLT is also a well-known dataset suitable for the evaluation of multi-session mapping [67]. Multiple sessions are collected from a mobile platform equipped with a Velodyne HDL-32E. Due to the rotational motion of the platform, we use FAST-LIO2 [4] for point cloud undistortion and LiDAR odometry generation. We selected 6 regions as our benchmark, with each region containing between 25 and 60 sessions, meaning the robot travels to the location more than 25 times.
@@
As mentioned above, we partition the large-scale urban areas into regions for evaluation. Essentially, the region represents a type of submap for long-term management, which differs from the concept of session. More specifically, the former is in the spatial domain, while the latter is in the temporal domain. We present the visualization results of the multi-session data and the regions in Figure 9 for a better illustration. Figure 10 also shows the mapping results in different regions. For visualization, a plane landmark is represented by a rhombus formed by selecting four points. These four control points are determined by computing the PCA of the weighted sum of the observed points within each landmark.
@@ VI-B Mapping Accuracy
LONER [^1] LONER fails to run on certain KITTI sequences, resulting in large drifts. We use $\times$ to indicate these “Failed” sequences. As stated in [54], LONER needs further revisions to operate city-scale scenarios. [54] 14.13 $\times$ 69.68 5.19 1.22 6.65 1.00 $\times$ 17.86 6.88 7.57 NA
PIN-LO [19] 5.6 4.3 9.3 0.7 0.1 1.7 0.5 0.5 3.0 1.8 0.8 2.6
PIN-SLAM [19] 0.8 4.3 2.1 0.7 0.1 0.3 0.4 0.3 2.1 1.2 0.8 1.2
@@
((a)) NCLT Region 04
((b)) HeLiPR DCC Region 05
((c)) HeLiPR KAIST Region 00
((d)) HeLiPR Town Region 05
This section validates that SLIM can provide a globally consistent map from multi-session data. We first compare SLIM with other advanced mapping systems on KITTI, using trajectory accuracy for quantitative evaluation. Then, we demonstrate long-term map merging and consistent mapping with incremental map sessions on the HeLiPR dataset.
@@ VI-B1 Mapping Accuracy on KITTI Dataset
Regarding the evaluation metrics, we compute the absolute rotation and translation error between the keyframes in our global map and the corresponding frames in the ground truth trajectory. The benchmark results are summarized in Table I, with partial results from BALM2 [12]. The overall results indicate that the advanced PIN-SLAM [19] achieves the best performance, while BALM2 [12] and proposed SLIM achieve competitive performance on certain sequences. We observe that SLIM performs poorly on Sequences 01 and 09 due to insufficient roadside features, such as streetlights and trees.
Among these comparisons, only the proposed SLIM, BALM2 [12], and PLC-LiSLAM [47] achieve feature-level LiDAR BA. BALM2 primarily focuses on state estimation rather than long-term mapping, as it lacks an explicit representation of landmarks. Additionally, BALM2 uses dense points for data association and residual construction and relies on a second-order solver. In contrast, PLC-LiSLAM uses planes, lines, and cylinders as landmarks with explicit representation. However, its landmark representation is not minimal, which necessitates additional design efforts to ensure correct state updates and maintain the sparsity of the nonlinear optimization problem. Our proposed SLIM employs a minimal landmark representation, ensuring that the corresponding optimization naturally exhibits a sparse structure. Furthermore, we design and validate the map merge and marginalization capabilities within the SLIM system, enhancing its overall functionality.
@@ VI-B2 Incremental Mapping on HeLiPR Dataset
Figure 11: Quantitative results of mapping accuracy on HeLiPR. The evaluation metric is the RMSE of absolute trajectory error. The overall results indicate that: (1) PGO (purple) and BA (cyan) effectively reduce errors and improve the consistency of the global map. (2) In certain cases, inaccurate submaps degrade the refinement results of PGO, while BA can almost entirely mitigate such impacts, demonstrating the necessity and effectiveness of our proposed BA.
Compared to the KITTI dataset, the recently published HeLiPR dataset is more suitable for validating long-term mapping. This is due to the inter-LiDAR sequences in HeLiPR, which naturally verify map merging and refinement as sessions increase at the same location. Specifically, the HeLiPR dataset encompasses sequences collected in four different scenarios: DCC, KAIST, Roundabout, and Town, with three sequences in each scenario. To obtain multiple submaps with overlapping regions, we selected 32 regions as our benchmark, with each region containing around 10 sessions. The SLIM system sequentially merges multiple submaps onto the base map, achieving incremental mapping from multi-session data. As depicted in Algorithm 1, SLIM conducts PGO and BA when a new session is merged, allowing for evaluation of the merged trajectory accuracy. We also compute the trajectory accuracy of the current submap to be merged. The heights of ground truth trajectories are not stable in some sequences of HeLiPR, so we evaluate trajectory accuracy along the $x$ and $y$ only.
((a)) Multi-session map on M2DGR
((b)) Map view at the campus gate
Figure 12: Demonstration in a university campus on M2DGR. For comparison, the point cloud maps are built by accumulating LiDAR points based on the robot poses estimated by SLIM.
((a)) DCC Region 06
((b)) Roundabout Region 00
Figure 13: Visualizing the BA effectiveness in HeLiPR. The point cloud maps are built by accumulating LiDAR points on robot poses provided by SLIM. The poses on the left side are obtained before the LiDAR BA, i.e., only using the PGO. After applying the BA, the poses on the right have fewer drifts, resulting in high-precision maps in zoomed-in views. This figure demonstrates the effectiveness of the proposed LiDAR BA on parameterized representations.
We present the experimental results of 20 sequences quantitatively in Figure 11. The SLIM system can maintain the consistency of the global map as new maps are added, even if the new maps are aligned with poor prior poses. Notably, the error of BA is generally less than PGO and the original LiDAR odometry, demonstrating the effectiveness of the proposed map refinement in Section IV-C. We also present qualitative results by applying our proposed BA in Figure 13, verifying the necessity of using BA for high-consistency mapping. We observe unstable results after applying PGO due to low-quality relative poses from block registration when odometry poses are not accurate enough. In summary, given multi-session maps, the map refinement modules can effectively reduce pose errors, i.e., improve mapping accuracy, by utilizing our proposed map merging and refinement.
One key capability of our framework is map merging, as described in Section IV-B, which is harder for quantitative evaluation compared to mapping accuracy. All the input submaps in this section are within their own coordinates when generated. The map merge module can align them into one global frame, resulting in the error reduction shown in Figure 11. The video in the supplementary materials also presents the map merge process to facilitate understanding.
In addition to the KITTI, NCLT and HeLiPR, we also test the SLIM system qualitatively on the M2DGR dataset [95].We present the qualitative mapping result in Figure 12. The SLIM system achieves map merging and refinement on the multi-session data provided by M2DGR, demonstrating its effectiveness in structural environments.
VI-C Memory Efficiency
TABLE II: Memory Consumption (Mega Byte, MB) on KITTI
Sequence (Length) 00 (3.7 Km) 05 (2.2 Km) 08 (3.2 Km)
Raw point cloud 13624.2 8284.7 12214.1
Point cloud ($r$=0.1m) 2831.46 1852.96 2769.22
Point cloud ($r$=0.3m) 52.03 32.58 70.41
Surfel map [96] 887.7 512.6 835.7
Mesh map [97] 2032.9 1317.4 1894.1
VDB TSDF map [98] 748.1 434.6 958.6
SHINE map [53] 160.6 114.2 189.9
LONER [^2] LONER has the same memory consumption across the sequences due to its map structure. The adaption of LONER to KITTI is available at https://github.com/qiaozhijian/LONER-KITTI. [54] 150.1 150.1 150.1
NeRF-LOAM [^3] Due to the limitations of GPU memory, NeRF-LOAM is performed on segments of KITTI sequences. [55] 953.61 800.69 1424.9
PIN-SLAM [19] 102.1 66.3 138.8
SLIM 12.9 7.2 9.2
SLIM (L) 0.5 0.3 0.4
((a)) NCLT Region 00
((b)) NCLT Region 01
((c)) NCLT Region 02
((d)) NCLT Region 03
Figure 14: Quantitative results of map consumption of SLIM and the time cost for map refinement. The green line represents “BA+Marginalization Time”; the pink line is “Landmark Consumption”; and the blue line is “Full Map Consumption”. The results on NCLT indicate that (1) the provided map is lightweight and only uses low storage consumption; (2) the consumption and time cost remain stable as the number of sessions increases.
Urban environments cover large spaces and various scenarios, so memory efficiency is a major concern for mobile robots. In this section, we evaluate the memory efficiency of SLIM compared to several conventional mapping systems in terms of storage size (memory consumption). The SLIM system utilizes lines and planes as basic front-end representations. The stored map can also contain graph structures that encode poses and co-visible information. If the map is solely designed for localization without such structures, we denote it with an (L) label for clarity. The first seven methods (from Eigen-Factor [17] to BALM2 [12]) in Table I utilize point clouds as explicit representation, resulting in dense point clouds as generated output maps.
The memory consumption of different maps is first evaluated on the KITTI dataset. The comparisons include both explicit representations, such as sampled point clouds (with $r$ as the sampling radius), and recent implicit neural representations [54, 53, 19, 55]. Partial results are from previous studies [8, 19], and all are summarized in Table II. Our lightweight map requires significantly less storage than other mapping systems, demonstrating its superiority in memory efficiency for urban environments. Specifically, SLIM (L) only needs less than 0.5 Megabyte (MB) to cover a travel length of over 3 kilometers (km), with a density of approximately 130 KB/km. If the full SLIM map is needed for updates, the system uses less than 15 MB per sequence to store on disks.
The memory consumption on NCLT and HeLiPR is presented in Figure 14 and Figure 15, respectively. In Figure 14, each region requires less than 1 MB to store a full map structure. Similar results are found in Figure 15. In summary, the built maps require low storage consumption on these two datasets, consistent with the maps generated from KITTI.
It is worth noting that the “efficiency” and “lightweight” in this study differ from those in recent LiDAR odometry [4, 99]. The two odometry systems focus on short-term and computationally efficient LiDAR mapping, while the SLIM is designed for long-term and memory-efficient mapping. High memory efficiency could reduce data transmission load in multi-robot systems for distributed operations, such as crowd-sourced urban mapping.
VI-D Scalability for Long-term Mapping
Sections VI-B and VI-C demonstrate that the SLIM system provides not only accurate but also memory-efficient maps in urban environments. The map merge capability is also verified in Section VI-B. As discussed in Section I, one motivation for SLIM is scalability, which is crucial for long-term mapping to maintain global consistency and bound map sizes as map sessions increase. Scalability is particularly important to ensure that map refinement, especially for high-dimensional LiDAR BA, remains controllable.
We verify the scalability of SLIM on NCLT and HeLiPR. For NCLT, Figure 14 shows that both map consumption and time cost remain within certain limits even with 50 sessions in the same region. This satisfying performance is primarily attributed to the marginalization module. To verify this, we disable the marginalization and conduct tests on HeLiPR. In Figure 15, if marginalization is disabled, we observe that map optimization time exhibits near-exponential growth, while map storage consumption increases linearly with the number of sessions. Overall, the full system maintains controllable costs on both NCLT and HeLiPR, verifying the scalability of SLIM. We do not provide experimental results on NCLT, as disabling marginalization will extend submap merging time to tens of minutes or even hours.
The proposed marginalization consists of two steps: topological reconstruction (Section V-B1) followed by map-centric NFR (Section V-B2), and the latter is one of the key contributions of this study. One might argue that only topological reconstruction or other advanced keyframe sampling techniques could maintain a sparse map structure for long-term mapping, making the NFR unnecessary due to its time costs in solving Equation (18). To verify this, we applied topological reconstruction (keyframe sparsification) in our method while maintaining the prior information described in Section IV-A1, i.e., without solving the NFR problem. Specifically, we define this prior information as $\sqrt{\mathbf{\Lambda}_{r_{k}}}=\mathbf{I}\cdot\sqrt{N_{r_{k}}/m}/\sigma$, considering the sum of points in the original observations, as the number of points directly reflects their weights. Figure 16 indicates that mapping accuracy drops if using priors compared to the full module (NFR enabled). The results demonstrate that NFR plays a vital role in both consistency and scalability for long-term mapping.
The minimum spanning tree generates the sparse topology of keyframe connections. Figure 17 shows the effectiveness of our keyframe sparsification approach, ensuring comprehensive connectivity among all keyframes while reducing unnecessary information storage. An additional issue in topology reconstruction is whether the system can preserve mapping accuracy when a landmark is retained with only a single observation on the keyframe, i.e., each landmark is connected only to the nearest keyframe pose, as described in Section V-B. Our prior experiments in Figure 11 show that the system consistently maintained high accuracy following each BA. The optimized results of the LiDAR BA remain unaffected by the precision of the multi-session input submaps. These results indicate that the proposed map-centric NFR can provide a reasonable information matrix to ensure the stability of mapping refinement, even when each landmark is connected to one keyframe pose after topology reconstruction.
((a)) Roundabout Region 00
((b)) Town Region 00
((c)) DCC Region 00
((d)) KAIST Region 00